OpenStudy (anonymous):

solve this equation if y=88 y=(0.48x)/(100.5-x)

OpenStudy (anonymous):

88=(0.48x)/(100.5-x) 88(100.5-x) = 0.48x 88(100.5) = 0.48x + 88x = 88.48x x= 88(100.5)/88.48, then use your calculator

OpenStudy (anonymous):

\[\text{Solve}\left[88\text{==} \left(\frac{12}{25}x\right)/\left(\frac{201}{2} - x\right),x\right]=\frac{55275}{553} \]

OpenStudy (anonymous):

Both solutions are 99.9548